i need help turning 2x+3y=9 into slope intercept form can you also provide an explanation for turning any standard form equation into slope intercept form
Answer by richwmiller(17219)

y=mx+b is the slope intercept form where m is the slope and b is the y intercept
so get y on one aside and everything else on the other by subtracting and dividing.
2x+3y=9
3y=-2x+9
y=-2x/3+9/3
y=-2/3*x+3